# MIFNet: Learning Modality-Invariant Features for Generalizable Multimodal Image Matching

### v1

@@ IV-E Ablation Study
Hyperparameters of Latent Diffusion Features. We analyse two hyper-parameters: 1) the layer $l$ of the decoder module in Stable Diffusion v2 [29], and 2) the steps $t$ of forward noise. As depicted in Fig. 6, the best result is achieved with the setting ($l=2,t=40$). This is probably because the features from layer $l=2$ provide better compromise between semantic and location accuracy. The step $t$ does not significantly influences the performance for a wide range of values. Using too large $t$ slightly decreases the performance.
@@
Furthermore, we observed that adding the LFA module without using the GMM clustering to optimize feature distribution did not lead to significant improvements. This is because the latent features output by stable diffusion contain noise, which hinders modality consistency after feature fusion. However, by incorporating the GMM clustering, the semantic distribution of the latent features becomes smoother, resulting in performance improvements of 7.5%, 5.4%, and 1.6% on the three datasets.
@@
Figure 9: Runtime analysis. We computed the average SRR and runtime of various methods on CF-OCT, CF-FA, and EMA-OCTA. Red: Methods that do not require cross-modal data for training. Blue: Methods that utilize data from different modalities for training.
IV-F Runtime Analysis
@@
We conduct experiments on a system equipped with an NVIDIA 3090 GPU (24 GB) and an 8-core Intel(R) Xeon(R) Platinum 8375C CPU. The average registration success rate and runtime are reported across three retinal sets. Fig. 9 presents the results of various methods. Among these, UR-SIFT-PIFD and SURF-PIFD-RPM are traditional algorithms that do not use GPU resources. Content-Adaptive and KPVSA-NET are multimodal methods specifically designed for fundus image matching, relying on modality transfer networks to reduce modality differences before matching, which limits their generalization ability. In contrast, MIFNet requires only training on color fundus images and performs zero-shot testing on unseen multimodal image pairs. Our method outperforms others in both accuracy and runtime. Specifically, MIFNet combined with SuperRetina achieves the highest success rate, while significantly outperforming traditional algorithms in speed, making it more suitable for clinical deployment.
@@ V Discussions
To better understand how the MIFNet improve the modality invariance and distinctiveness of the features, we conduct two analysis in this section. Without losing generality, we employ XFeat [20] as the feature extractor in our analysis.
V-A Modality Invariance Analysis

### v3

@@ IV-E Ablation Study
Hyperparameters of Latent Diffusion Features. We analyse two hyper-parameters: 1) the layer $l$ of the decoder module in Stable Diffusion v2 [29], and 2) the steps $t$ of forward noise. As depicted in Fig. 6, the best result is achieved with the setting ($l=2,t=40$). We observe that features from decoder layer $l=2$ consistently outperform those from shallower ($l=1$) or deeper layers ($l=3$, $l=4$), indicating that intermediate layers better balance semantic abstraction and spatial detail. Additionally, the performance remains stable across a wide range of time steps $t$, demonstrating robustness to latent noise, though a slight drop occurs when $t>100$ due to weakened semantic coherence. These findings are crucial for designing effective and robust modality-invariant feature learning.
@@
Furthermore, we observed that adding the LFA module without using the GMM clustering to optimize feature distribution did not lead to significant improvements. This is because the latent features output by stable diffusion contain noise, which hinders modality consistency after feature fusion. However, by incorporating the GMM clustering, the semantic distribution of the latent features becomes smoother, resulting in performance improvements of 7.5%, 5.4%, and 1.6% on the three datasets. We also experimented with K-means clustering as an alternative, but it performed worse than GMM due to its inability to model soft assignments and complex feature distributions.
@@
(a) Number of attention layers $L$
(b) Foundation model type
Figure 9: Ablation study of MIFNet on (a) the number of attention layers and (b) different foundation models (DINOv2 vs. Stable Diffusion v2) using XFeat as the base matcher.
Influence of the Number of Attention Layers. We evaluate the impact of the number of attention layers in MIFNet for $L\in\{1,3,5,7,9,11\}$ and report the average registration success rate (SRR) on three fundus datasets: CF-OCT, CF-FA, and EMA-OCTA. XFeat [20] is used as the base feature extractor. As shown in Fig. 9(a), increasing $L$ from 1 to 9 improves performance. The results with $L=9$ and $L=11$ are comparable. We choose $L=9$ as the default setting to maintain a good balance between accuracy and inference time.
Influence of Different Foundation Models. We compare two widely-used pretrained foundation models, Stable Diffusion v2 (SDv2) and DINOv2, to investigate their impact on MIFNet’s cross-modal matching performance.
As shown in Fig. 9(b), both models improve upon the original XFeat baseline across all datasets, with SDv2 consistently outperforming DINOv2. In the fundus imaging scenario, SDv2 achieves higher SRR than DINOv2 by 11.5%, 18.3%, and 20.0% on the CF-OCT, CF-FA, and EMA-OCTA datasets, respectively. The performance gap is more pronounced in datasets with greater modality discrepancies. In the remote sensing scenario, the improvement margins are relatively smaller, with gains of 4.4% on Optical-SAR and 2.2% on Optical-NIR, likely due to the higher visual similarity between remote sensing images and the natural images used during pretraining. Compared to DINOv2, SDv2 produces more modality-invariant features in zero-shot settings, owing to its stronger semantic representation capability. These results suggest that integrating even more powerful foundation models could further improve the overall performance of MIFNet.
TABLE IV: Influence of training on different modalities for multimodal image matching. MIFNet(O): trained on optical images, MIFNet(S): trained on SAR images, MIFNet(O+S): trained on a mixed dataset composed of half optical and half SAR images.
Methods Optical - SAR Optical - NIR
[…]
XFeat 18.4 3.0 77.4 37.8
MIFNet (O) 36.1 8.1 85.9 46.1
MIFNet (S) 39.3 9.2 83.0 45.3
MIFNet (O+S) 37.5 8.7 85.5 45.9
Influence of Training Dataset Composition. Our approach is trained using synthetic image pairs generated by applying random geometric transformations to individual images, without requiring any paired multimodal data. To assess the effect of training data composition, we conduct experiments with three configurations: optical only (O), SAR only (S), and a mixed set of unpaired optical and SAR images (O+S). Each single-modality dataset contains 2,400 images, while the mixed set includes 1,200 images from each modality.
As shown in Tab. IV, all three training configurations outperform the baseline (XFeat). In Optical-SAR scenario, SAR-only training achieves the best SRR (39.3%). The mixed dataset also improves performance by 1.4% over Optical-only. In contrast, in Optical-NIR scenario, SAR-only training yields the lowest performance (83.0%). The mixed dataset slightly underperforms compared to optical-only. These results suggest that incorporating data from modalities similar to the test domain can improve performance. Nevertheless, since optical data are widely available and already effective, we adopt optical-only training in all other experiments.
Figure 10: Average successful registration rate (SRR) across three fundus datasets (CF-OCT, CF-FA, EMA-OCTA) under different RMSE thresholds {1, 3, 5, 10} pixels.
Effect of Varying RMSE Thresholds. To further assess the robustness of our MIFNet, we report the successful registration rate (SRR) under three commonly used RMSE thresholds: 1, 3, 5, and 10 pixels. As shown in Fig. 10, our proposed MIFNet improves the robustness of baseline detectors such as SuperPoint [21], SuperRetina [36], and XFeat [20] across all thresholds. Moreover, when combined with SuperRetina, MIFNet consistently outperforms other baselines, including Xoftr [62] and GIM [66], under different thresholds.
IV-F Runtime Analysis and Computational Complexity
@@
Figure 11: Runtime analysis. We computed the average SRR and runtime of various methods on CF-OCT, CF-FA, and EMA-OCTA. Red: Methods that do not require cross-modal data for training. Blue: Methods that utilize data from different modalities for training. The number inside each circle represents the computational cost in FLOPs.
@@
We conduct experiments on a system equipped with an NVIDIA 3090 GPU (24 GB) and an 8-core Intel(R) Xeon(R) Platinum 8375C CPU. The runtime and computational complexity are important metrics for applying image matching to downstream tasks. The average successful rate and runtime are reported across three retinal datasets. Both runtime and GFLOPs are measured as the total latency and computational cost for registering one image pair. Fig. 11 presents the results of various methods. Among these, UR-SIFT-PIFD [42] and SURF-PIFD-RPM [18] are traditional algorithms that do not use GPU resources. Content-Adaptive [45], KPVSA-NET [25], and Xoftr [62] are learning-based multimodal methods that rely on paired cross-modal training data. In contrast, MIFNet requires only training on color fundus images and performs zero-shot testing on unseen multimodal image pairs. Specifically, MIFNet combined with SuperRetina achieves the highest success rate and significantly outperforms traditional algorithms in terms of speed, while being only marginally slower than Xoftr. The main computational cost comes from Stable Diffusion v2, which is used during inference for latent feature extraction. SuperRetina + SDv2 + MIFNet requires a total of 1856.2 GFLOPs, with SuperRetina accounting for 301.2 GFLOPs and SDv2 for 1522.7 GFLOPs. In comparison, MIFNet’s network layers only add 32.3 GFLOPs and 11.3M parameters, making it lightweight and efficient.
@@ V Discussions
To better understand how MIFNet improves the geometric invariance, modality invariance, distinctiveness, and generality of the learned features, we conduct four analyses in this section. Without losing generality, we employ XFeat [20] as the feature extractor.
TABLE V: Successful registration rate under rotation perturbations.
Methods CF-OCT CF-FA EMA-OCTA Opt-SAR Opt-NIR
Xoftr 48.1 41.3 6.7 19.8 45.0
GIM 73.1 36.5 0 6.8 48.5
SuperPoint 57.7 8.8 0 2.5 28.5
SP+MIFNet 86.5 60.0 13.3 20.0 53.5
ALIKED 15.4 6.6 0 1.3 30.5
AL+MIFNet 69.2 48.8 26.7 18.5 55.0
XFeat 17.3 3.5 0 4.0 31.0
XF+MIFNet 88.5 57.5 46.7 28.3 68.5
V-A Geometric Invariance Analysis
To evaluate the robustness of our method under rotational variations, we apply random rotations to the input image pairs: fundus images are rotated within $[-36^{\circ},36^{\circ}]$, and remote sensing images within $[-90^{\circ},90^{\circ}]$. As shown in Tab. V, the performance of lightweight single-modality detectors such as SuperPoint, ALIKED, and XFeat degrades significantly under increased rotation. In contrast, our method exhibits strong robustness to such geometric transformations. Notably, when combined with XFeat, our method achieves the best performance across all test datasets under rotation conditions.
V-B Modality Invariance Analysis
@@ V-C Distinctiveness Analysis
TABLE VI: Generalization results on the FIRE dataset (color fundus–color fundus) and the SRIF dataset (optical–optical). The MIFNet models used here are trained on the FA modality and the SAR modality, respectively.
Methods FIRE (CF–CF) SRIF (Opt–Opt)
SRR(%) $\uparrow$ RMSE $\downarrow$ SRR(%) $\uparrow$ $H_{\text{err}}$downarrow
ALIKED 85.0 13.4 32.5 3.3
ALIKED + MIFNet 95.1 5.1 48.5 2.6
XFeat 91.7 8.0 38.0 3.1
XFeat + MIFNet 97.7 3.9 52.5 2.5
V-D Generalization Capability Across Different Modalities
To evaluate the cross-modal generalization capability of our MIFNet, we conduct additional experiments where the model is trained on one modality and tested on image pairs from another different modality. In the fundus imaging scenario, MIFNet is trained using 400 fluorescein angiography (FA) images and evaluated on the FIRE dataset [69], which contains 134 color fundus (Color Fundus–Color Fundus) image pairs. In the remote sensing scenario, the model is trained on 2,400 SAR images and tested on Optical–Optical image pairs from the SRIF dataset [70].
As shown in Tab. VI, even though both the training and test data for these baselines come from similar modalities, our method still improves the performance of baseline methods (ALIKED [41], XFeat [20]). This highlights the strong generalization ability of MIFNet, which effectively improves the matching performance of baseline models across different modalities.
Figure 14: Visualization of some failure cases.
V-E Failure Case Analysis
Although our method achieves strong performance across various scenarios, it still faces challenges in certain cases. As shown in Fig. 14, on the left, large cross-modal appearance gaps lead to poor keypoint repeatability, which limits our method since it mainly enhances descriptor robustness. On the right, heavy noise and lack of semantic structure (e.g., in degraded SAR images) also affect matching accuracy. Addressing these issues will be part of our future work.
@@ VI Conclusion
VII Acknowledgment
This work was supported in part by the National Key Research and Development Program of China (Grant No. 2023YFC2705700), the National Natural Science Foundation of China (Grant Nos. 62222112, 62225113, and 62176186), the Innovative Research Group Project of Hubei Province (Grant No. 2024AFA017), and the Key Research and Development Program of Zhejiang Province (Grant No. 2024C03204). This project was also supported by A*STAR under the Coastal Protection and Flood Management Research Programme (Grant No. 2024/T2/H1-P4).